OpenStudy (anonymous):

Please Please help! Would appreciate it alot! How many solutions are there to the following system of equations?

OpenStudy (anonymous):

6x – 9y = –9 and

OpenStudy (anonymous):

A. 1 B. 2 C. infinitely many D. 0

hartnn (hartnn):

plug in the value of y from 2nd equation into 1st equation that is put y =2/3 x+1 in 6x-9y=-9

OpenStudy (anonymous):

just 2, because the second one seems like its for y, and there's only x and y we have to solve for, so 2

OpenStudy (anonymous):

hmmm....so @hartnn are you saying 2 or not not

hartnn (hartnn):

no, there are more than 2 how ? we need to find out so did u do what i asked for ?

OpenStudy (anonymous):

well 6x -9(2/3x+1) = -9 6x -6x -9 = -9 0 = 0

OpenStudy (anonymous):

yes i plugged it in

OpenStudy (anonymous):

now what

hartnn (hartnn):

what did u get ?

OpenStudy (anonymous):

so its 2

hartnn (hartnn):

@jvlinh so which 2 solutions you got ?

OpenStudy (anonymous):

6x-9(y =2/3 x+1)=-9

OpenStudy (anonymous):

is what i got

OpenStudy (anonymous):

only x and y are in there, and the second only has y, so 2?

OpenStudy (anonymous):

is it true @hartnn

hartnn (hartnn):

thats not the way to solve and no

hartnn (hartnn):

" 6x -9(2/3x+1) = -9 6x -6x -9 = -9 0 = 0" did u get that process^ ??

hartnn (hartnn):

we got 0=0, right ? is it true for any values of x and y ?? or for only some values ?

hartnn (hartnn):

you still here ? @TakeA.I.M. ?

OpenStudy (anonymous):

yes sry i was gone for a min

OpenStudy (anonymous):

its true for values x and y right?

hartnn (hartnn):

yes, for all or any values of x and y thats why we have infinite solutions :)

OpenStudy (anonymous):

ok thnx

hartnn (hartnn):

welcome ^_^